OpenStudy (anonymous):

Find all Pythagorean triples with 3 sides in in arithmetic progression?

OpenStudy (anonymous):

I know what a Pythagorean triple is, but what do you mean by one having three sides?

OpenStudy (anonymous):

Right triangle with Pythagorean triple is a Pythagorean triangle

OpenStudy (anonymous):

Think of it as a bit of a hint....

OpenStudy (anonymous):

3,4,5 the basic thing!! let me think about the rest

OpenStudy (anonymous):

Ofcourse by 3,4,5 I also mean all such triplets..6,8,10 ; 9,12,15 etc

OpenStudy (anonymous):

I think there are no more except the 3-4-5 forms

OpenStudy (anonymous):

Thinking is good, a proof would be better...:-)

OpenStudy (anonymous):

Hint?

OpenStudy (anonymous):

(a+d)^2 = a^2 + (a-d)^2 Solving, 4d = a Putting different values of d, we get different values of a..which are all multiples of 4. a-d are respective multiples of 3 a+d are respective multples of 5

OpenStudy (anonymous):

That will do....

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

x^2 + (x+d)^2 = (x+2d)^2 -> x^2 -2xd -3d^2 = 0 -> (x-3d)(x+d) = 0 etc.....